Consider the situation of the previous problem. (a) Find the tension in the string in equilibrium. (b) Suppose the ball is slightly pushed aside and released. Find the time period of the small oscillations.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
(a) Tension in the string in Equilibrium T cos 60° = mg ⇒ T = mg/cos 60° = 10 * 10^-3 * 10/1/2 = 10^-1 * 2 = 0.20 N (b) Straingtening the same figure Now the resultant for ‘R’ Induces the acceleration in the pendulum. T = 2 * π √ℓ/g = 2π √ℓ/[g^2 +(σq/2ε base 0m)^2]^1/2 = 2π √ℓ/[100 +(0.2 * √3/2 * 10^-2)^2]^1/2 = 2π √ℓ/(100 + 300)^1/2 = 2π √ℓ/20 = 2 * 3.1416 * √10 * 10^-2/20 = 0.445 sec.